# Graham’s rearrangement for dihedral groups

### v1

@@
Abstract
A famous conjecture of Graham asserts that every set $A\subseteq\mathbb{F}_{p}\setminus\{0\}$ can be ordered so that all partial sums are distinct. Bedert and Kravitz proved in [4] that this statement holds whenever $|A|\leqslant e^{c(\log p)^{1/4}}$.
In this paper, we will use a similar procedure to obtain an upper bound of the same type in the case of dihedral groups $Dih_{p}$.
1 Introduction
Let $A$ be a finite subset of group $(G,\cdot)$. We say that an ordering $a_{1},\ldots,a_{|A|}$ of $A$ is valid if the partial products (or partial sums in additive notation) $p_{1}=a_{1},p_{2}=a_{1}\cdot a_{2},\ldots,p_{|A|}=a_{1}\cdot a_{2}\cdot\cdots\cdot a_{|A|}$ are all distinct. Moreover, this ordering is a sequencing if it is valid and $p_{i}\not=id$ for any $i\in[1,|A|-1]$. In this case, we will say that $A$ is sequenceable. In the literature, there are several conjectures about valid orderings and sequenceability. We refer to [8, 12, 14] for an overview of the topic, [1, 2, 3, 7] for lists of related conjectures and [5]. Here, we explicitly recall just that of Graham, who conjectured that every set of non-zero elements of $\mathbb{F}_{p}$ has a valid ordering.
@@ 1 Introduction
Let $p$ be a prime. Then every subset $A\subseteq\mathbb{F}_{p}\setminus\{0\}$ has a valid ordering.
Until recently, the main results on this conjecture were for small values of $|A|$; in particular, in [7], the conjecture was proved for sets $A$ of size at most $12$.
The first result involving arbitrarily large sets $A$ was presented last year by Kravitz [11]. He used a simple rectification argument to show that Graham’s conjecture holds for all sets $A$ of size $|A|\leqslant\log p/\log\log p$. A similar argument was also proposed (but not published) by Will Sawin [15] in a 2015 MathOverflow post.
@@
Let $p$ be a large enough prime and let $c>0$. Then every subset $A\subseteq\mathbb{F}_{p}\setminus\{0\}$ is sequenceable provided that
$|A|\leqslant e^{c(\log p)^{1/4}}.$
Graham’s conjecture over non-abelian groups was first studied by Ollis in [13] over dihedral groups. More recently, the authors in [5] presented a nice connection with rainbow paths and produced the asymptotic result that a subset $A$ of a finite group $G$ admits an ordering in which at least $(1-o(1))|A|$ many partial products are distinct. Returning to dihedral groups, inspired by the result of Kravitz ([11]) on cyclic groups, Costa and Della Fiore proved in [6] that any subset $A$ of the dihedral group $Dih_{p}\setminus\{id\}$ (and in general, of a class of semidirect products) is sequenceable, provided that $p>\frac{|A|!}{2}$. Note that this is slightly better than $|A|\leq\log p/\log\log p$. This paper is dedicated to proving, with a similar procedure as that of Bedert and Kravitz, that subsets of the dihedral group $Dih_{p}\setminus\{id\}$ are sequenceable if they are not too large with respect to $p$. In this setting, we improve the results of [6] on dihedral groups, obtaining essentially the same bound as in the case of $\mathbb{F}_{p}$. That is,
@@
Let $p$ be a large enough prime and let $c>0$. Then every subset $A\subseteq Dih_{p}\setminus\{id\}$ is sequenceable provided that
$|A|\leqslant e^{c(\log p)^{1/4}}.$
2 Proof Strategy
@@
We offer a brief summary of the main steps of our proof of Theorem 1.3. Note that we represent elements in $Dih_{p}$ as $(x_{i},a_{i})$ where $x_{i}\in\mathbb{F}_{p}$ and $a_{i}\in\mathbb{F}_{2}$.
@@
We begin Section 3 by describing how a small set $B\subseteq Dih_{p}$ can be “rectified” such that all of its $\mathbb{F}_{p}$ components are contained within a small interval. Then, for a nonempty subset $A\subseteq Dih_{p}\setminus\{id\}$, we prove the existence of an automorphism $\phi\in Aut(Dih_{p})$ such that $\phi(A)=E\cup\left(\cup_{j=1}^{s}D_{j}\right)$, where each $D_{j}$ is a dissociated set with size a multiple of eight and where the $\mathbb{F}_{p}$ components of $E$ are in a small interval by our rectification argument. Moreover, we can impose that each dissociated set $D_{j}$ has $\mathbb{F}_{2}$ components either constantly-one, (in this case we say that $j\in L_{1}$) or constantly 0 ($j\in L_{0}$). Here we split each $D_{j}$ with $j\in L_{1}$ into an odd-position subset $D^{o}_{j}$ and an even-position subset $D^{e}_{j}$ of about equal size. Then the product $\tau_{j}$ of its elements, with elements of $D^{o}_{j}$ in odd position and elements of $D^{e}_{j}$ in even position, does not depend on the ordering of those elements. This means that, even if $Dih_{p}$ is non-abelian, we have a constant product $\delta$ of the elements of the dissociated sets without fixing the ordering of all the elements. We then relabel $\phi(A)$ by $A$ for ease of notation.
@@
In Section 4, we order the elements of $E$. Here we split $E$ into sets $P$ (with $\mathbb{F}_{p}$ components positive and $\mathbb{F}_{2}$ components constantly 0), $N$ (with $\mathbb{F}_{p}$ components negative and $\mathbb{F}_{2}$ components constantly-zero), and $S$ (with $\mathbb{F}_{2}$ components constantly-one). We further split $S$ into subsets $S_{e}$ and $S_{o}$ of almost equal size, where the $\mathbb{F}_{p}$ component of each element of $S_{e}$ is larger than the $\mathbb{F}_{p}$ components of all of the elements in $S_{o}$. We then prove the existence of an ordering $\mathbf{x}$ of the elements of $P,N,S_{e},$ and $S_{o}$ such that $\delta,\mathbf{x}$ is a sequencing, where $\delta$ is the prespecified product of the elements of the $D_{j}$s. At the same time, we derive an important bound on the size of overlap between subsets of $Dih_{p}$ and $\delta$ times the partial products of $\mathbf{x}$}.
@@
Finally, in Section 5 of our paper, we describe a two-step method of ordering the $D_{j}$s. First, we randomly split each $D^{o}_{j}$, each $D^{e}_{j}$ (when $j\in L_{1}$), and each remaining $D_{j}$ (when $j\in L_{0}$) into a four-set partition. Here, we use a probabilistic argument to prove that, given a sequencing $\mathbf{x}$ of $E$, we can obtain a sequencing $\tau_{1},\dots\tau_{u}\mathbf{x}$ of $E$ and the products $\tau_{i}$ of the partitions of the $D_{j}$s. To conclude, we use again a probabilistic argument to prove that if $\mathbf{x}$ satisfies a technical condition (on the intersection between suitable subsets and $\delta$ times the partial products of $\mathbf{x}$}), there exists a sequencing of $E$ and each partition $T_{i}$ of the $D_{j}$s. This gives us Theorem 1.3.
3 Structure Theorem
@@
In this paper we will consider groups that are semidirect products $G=\mathbb{F}_{p}\rtimes_{\varphi}H$ for some group homomorphism $\varphi:H\rightarrow Aut(\mathbb{F}_{p})$. We recall that the group operation is defined as
@@
is an automorphism of $(\mathbb{F}_{p},+)$. It is well known that the automorphisms of $(\mathbb{F}_{p},+)$ are the multiplications and hence that there exists $\varphi_{a_{1}}\in\mathbb{F}_{p}$ such that
@@
Here we let $Dih_{p}$ denote the dihedral group $\mathbb{F}_{p}\rtimes_{\varphi}\mathbb{F}_{2}$, where $\varphi(1)(x)=-x$ and $\varphi(0)(x)=x$. In this case, the group operation is defined as
@@
where $\varphi_{0}=1$ and $\varphi_{1}=-1$.
Given a non necessarily abelian group $G$, a subset $D=\{d_{1},\ldots,d_{r}\}\subseteq G$ is dissociated if
@@
Let $B\subseteq G$ be a finite subset of a non necessarily abelian group. If $D=\{d_{1},\ldots,d_{r}\}$ is a maximal dissociated subset of $B$, then
@@
$d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots d_{\sigma(i-1)}^{\epsilon(i-1)}\cdot b\cdot d_{\sigma(i)}^{\epsilon(i)}\cdots\cdot d_{\sigma(r)}^{\epsilon_{\sigma(r)}}=id.$
@@
$b=\left(d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots d_{\sigma(i-1)}^{\epsilon(i-1)}\right)^{-1}\left(d_{\sigma(i)}^{\epsilon(i)}\cdots\cdot d_{\sigma(r)}^{\epsilon_{\sigma(r)}}\right)^{-1}$
@@
Also, the following lemma relates the size of a set and its dimension.
@@
Let $B$ be a finite subset of $Dih_{p}$. Then $|B|\leqslant 3^{\dim(B)}$.
@@
Let $D={(x_{1},a_{1}),(x_{2},a_{2}),\dots,(x_{r},a_{r})}$ be a maximal dissociated subset of $B$ of size $dim(B)$. Due to the previous lemma, this means that any element $(x,a)$ of $B$ belongs to $Span(D)$ and hence
@@
$(x,a)$ $=\left(x_{\sigma(1)}+\varphi_{a_{\sigma(1)}}x_{\sigma(2)}+\dots+\varphi_{a_{\sigma(1)}+\cdots+a_{\sigma(t-1)}}x_{\sigma(t)},\sum_{i=1}^{t}a_{\sigma(i)}\right)$
@@
where $(\gamma_{i},\lambda_{i})\in\{(0,0),(1,1),(-1,1)\}$. Since we have three choices for each pair $(\gamma_{i},\lambda_{i})$, it follows that $|B|\leqslant 3^{\dim(B)}$. ∎
Following [4], we state a lemma that allows us to “rectify” small sets. We define, for each (nonempty) subset $A\subseteq Dih_{p}$, the parameter
$R=R(A)\vcentcolon=c_{1}\max\left((\log p)^{1/2},\frac{\log p}{\log|A|}\right),$ (1)
@@
If $B\subseteq Dih_{p}$ is a nonempty subset of dimension $\dim(B)<R=R(B)$, then there is some $\phi\in Aut(Dih_{p})$ such that the image $\pi_{1}(\phi(B))$ (where $\pi_{1}$ is the projection over the $\mathbb{F}_{p}$ component) is contained in the interval $(-\frac{p}{100|B|},\frac{p}{100|B|})$.
@@
Given $\lambda\in\mathbb{F}_{p}$, we consider here the family of maps $\phi_{\lambda}:Dih_{p}\rightarrow Dih_{p}$ defined as
@@
Note that, if $\lambda\not=0$, $\phi_{\lambda}\in Aut(Dih_{p})$.
@@
$\{\lambda(x_{1},\dots,x_{r}):\lambda\in\mathbb{F}_{p}\}\subseteq(\mathbb{F}_{p})^{r}.$
Due to the pigeonhole principle, there exists distinct $\lambda_{1},\lambda_{2}\in\mathbb{F}_{p}$ such that $\|\lambda_{1}x_{i}-\lambda_{2}x_{i}\|\leqslant p^{1-1/r}$ for all $i\in[1,r]$. Set $\lambda\vcentcolon=\lambda_{1}-\lambda_{2}$, we have that $\phi_{\lambda}\in Aut(Dih_{p})$ and $\pi_{1}(\phi_{\lambda}(D))\subseteq[-p^{1-1/\dim(B)},p^{1-1/\dim(B)}]$. Since $B\subseteq\operatorname{span}(D)$, we have
@@
The thesis follows because we can prove, as done in [4], that $\dim(B)p^{1-1/\dim(B)}<p/(100|B|)$ as long as $c_{1}$ is chosen to be sufficiently small. ∎
Before stating the structure theorem for dihedral groups, we prove the following Lemma, which allows us to modify a dissociated set.
@@
Let $G$ be a non necessarily abelian group, and let $D_{1}\cup D_{2}$ be a partition of a dissociated subset $D$ of $G$. For every element $x\in G\setminus\{0\}$, either $D_{1}\cup\{x\}$ or $D_{2}\cup\{x\}$ is dissociated.
@@
Now we show that any subset $A$ of $Dih_{p}$ can be split in a family of dissociated sets and a remainder $E$ of small dimension that can be embedded, as a consequence of the rectification lemma, into $\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)$. Also, to avoid the cases already covered by Theorem 1.2, we assume that $A$ is not contained in $\{(x,0):x\in\mathbb{F}_{p}\}$ which is isomorphic to $\mathbb{F}_{p}$.
@@
Let $p$ be large enough. For every nonempty subset $A\subseteq Dih_{p}\setminus\{id\}$ not contained in $\{(x,0):x\in\mathbb{F}_{p}\}$, there is some $\phi\in Aut(Dih_{p})$ such that $\phi(A)$ can be partitioned as
@@
    Each $D_{j}$ is a dissociated set of size $\Theta(R)$ a multiple of eight.
@@
    We can split the interval $[1,s]=L_{0}\cup L_{1}$ so that $1$ and $s$ belong to the same set $L_{i}$, for any $j\in L_{1}$, $D_{j}\subseteq\{(x,1):x\in\mathbb{F}_{p}\}$ and, for any $j\in L_{0}$, $D_{j}\subseteq\{(x,0):x\in\mathbb{F}_{p}\}$.
@@
    Given $j\in L_{1}$, we split $D_{j}$ into $D_{j}^{o}$ and $D_{j}^{e}$ so that $|D_{j}^{o}|=|D_{j}^{e}|=|D_{j}|/2$. Then we set
    $\delta\vcentcolon=\left(\prod_{j\in L_{1}}\prod_{d^{o}\in D_{j}^{o},d^{e}\in D_{j}^{e}}d^{o}\cdot d^{e}\right)\left(\prod_{j\in L_{0}}\prod_{d\in D_{j}}d\right)$
    (which does not depend on the orderings on $D_{j}^{o}$ and $D_{j}^{e}$). Then $\delta=(z_{0},0)$ where $z_{0}\not=0$ and

### v2

# Graham’s rearrangement for a class of semidirect products

@@
Abstract.
A famous conjecture of Graham asserts that every set $A\subseteq\mathbb{Z}_{p}\setminus\{0\}$ can be ordered so that all partial sums are distinct. Bedert and Kravitz proved in [4] that this statement holds whenever $|A|\leq e^{c(\log p)^{1/4}}$.
In this paper, we will use a similar procedure to obtain an upper bound of the same type in the case of semidirect products $\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $\varphi:H\to Aut(\mathbb{Z}_{p})$ satisfies $\varphi(h)\in\{id,-id\}$ for each $h\in H$ and where $H$ is abelian and each subset of $H$ can be ordered such that all of its partial products are distinct.
Key words and phrases:
Sequenceability, Rectification
1991 Mathematics Subject Classification:
11B75
1. Introduction
Let $A$ be a finite subset of group $(G,\cdot)$. We say that an ordering $a_{1},\ldots,a_{|A|}$ of $A$ is valid if the partial products (or partial sums in additive notation) $p_{1}=a_{1},p_{2}=a_{1}\cdot a_{2},\ldots,p_{|A|}=a_{1}\cdot a_{2}\cdot\cdots\cdot a_{|A|}$ are all distinct. Moreover, this ordering is a sequencing if it is valid and $p_{i}\not=id$ for any $i\in[1,|A|-1]$. In this case, we say that $A$ is sequenceable. We also say that a group $G$ is strongly sequenceable if all the subsets of $G\setminus\{id\}$ are sequenceable. In the literature, there are several conjectures about valid orderings and sequenceability. We refer to [8, 12, 14] for an overview of the topic, [1, 2, 3, 7] for lists of related conjectures, and [5] for a treatment using rainbow paths. Here, we explicitly recall only the conjecture of Graham.
@@ 1. Introduction
Let $p$ be a prime. Then every subset $A\subseteq\mathbb{Z}_{p}\setminus\{0\}$ has a valid ordering.
Until recently, the main results on this conjecture were for small values of $|A|$; in particular, the conjecture was proved for sets $A$ of size at most $12$ [7].
The first result involving arbitrarily large sets $A$ was presented last year by Kravitz [11]. He used a simple rectification argument to show that Graham’s conjecture holds for all sets $A$ of size at most $\log p/\log\log p$. A similar argument was also proposed (but not published) by Sawin [15] in a 2015 MathOverflow post.
@@
Let $c>0$ be a constant and $p$ be a large enough prime. Then every subset $A\subseteq\mathbb{Z}_{p}\setminus\{0\}$ is sequenceable provided that
$|A|\leq e^{c(\log p)^{1/4}}$
Graham’s conjecture over non-abelian groups was first studied by Ollis in [13] over dihedral groups. More recently, the authors in [5] presented a nice connection with rainbow paths and produced the asymptotic result that any subset $A$ of a finite group $G$ admits an ordering in which at least (1o(1))|A|(1-o(1))|A| many partial products are distinct. Inspired by the result of Kravitz ([11]) on cyclic groups, Costa and Della Fiore proved in [6] that any subset $A$ of $Dih_{p}\setminus\{id\}$, where $p$ is prime and $Dih_{p}$ is the dihedral group of size $2p$, and in general any subset $A$ of a class of semidirect products, is sequenceable provided that $p>|A|!/2$. Note that this is slightly better than $|A|\leq\log p/\log\log p$.
Section 3 of this paper is dedicated to proving, with a similar procedure, the sequenceability of subsets of semidirect products $\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $\varphi:H\to Aut(\mathbb{Z}_{p})$ satisfies $\varphi(h)\in\{id,-id\}$ for each $h\in H$ and where $H$ is abelian and strongly sequenceable. We note that for $H=\mathbb{Z}_{2}$, we are considering the dihedral group $Dih_{p}$. In this setting, we improve the results of [6] on dihedral groups, obtaining essentially the same bound as in the case of $\mathbb{Z}_{p}$ (with a different best possible constant $c$). That is,
@@
Let $H$ be a finite abelian and strongly sequenceable group. There exists a $c>0$ such that every subset $A\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H\setminus\{id\}$, where $\varphi:H\to Aut(\mathbb{Z}_{p})$ satisfies $\varphi(h)\in\{id,-id\}$ for each $h\in H$, of size
$|A|\leq e^{c(\log p)^{1/4}},$
has a sequencing.
The main difficulty in adapting the approach of [4] to get this theorem is that the semidirect products are not necessary abelian and hence the product of the elements of a set is not fixed. For this reason, we have to impose some restrictions on the positions of certain elements.
2. Proof Strategy
@@
We offer a brief summary of the main steps of our proof of Theorem 1.3. Note that we represent elements in $\mathbb{Z}_{p}\rtimes_{\varphi}H$ as $(x_{i},a_{i})$ where $x_{i}\in\mathbb{Z}_{p}$ and $a_{i}\in H$.
@@
We begin Section 3 by describing how a small set $B\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$ can be “rectified” such that all of its $\mathbb{Z}_{p}$ components are contained within a small interval. Then, for a nonempty subset $A\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H\setminus\{id\}$, we prove the existence of an automorphism $\phi\in Aut(\mathbb{Z}_{p}\rtimes_{\varphi}H)$ such that $\phi(A)=E\cup\left(\cup_{j=1}^{s}D_{j}\right)$, where each $D_{j}$ is a dissociated set with size a multiple of eight and where the $\mathbb{Z}_{p}$ components of $E$ are in a small interval by our rectification argument. Moreover, we can impose that each dissociated set $D_{j}$ has elements $d=(x,a)\in D_{j}$ such that either $\varphi(a)=-id$ for all $a$ (in which case we say that $j\in L_{1}$) or $\varphi(a)=id$ for all $a$ ($j\in L_{0}$). Here we split each $D_{j}$ with $j\in L_{1}$ into an odd-position subset $D^{o}_{j}$ and an even-position subset $D^{e}_{j}$ of about equal sizes. Then the product $\tau_{j}$ of the elements of $D_{j}$, with elements of $D^{o}_{j}$ in odd position and elements of $D^{e}_{j}$ in even position, does not depend on the orderings within $D_{j}^{o}$ and $D_{j}^{e}$. This means that, even if $\mathbb{Z}_{p}\rtimes_{\varphi}H$ is non-abelian, we have a constant product $\delta$ of the elements of the dissociated sets without fixing the ordering of all the elements. We then relabel $\phi(A)$ by AA for ease of notation.
@@
In Section 4, we order the elements of $E$. Here we split $E$ into sets $P$ (with $\mathbb{Z}_{p}$ components positive [^1] Here we say the positive elements are the ones in the interval $1,\dots,(p-1)/2]$ while the negatives in [(p1)/2,1][-(p-1)/2,-1]. and $H$ components such that $\varphi(a)=id$), NN (with $\mathbb{Z}_{p}$ components negative and $H$ components such that $\varphi(a)=id$), $Z$ (with $\mathbb{Z}_{p}$ components always zero and $H$ components such that $\varphi(a)=id$), and $S$ (with $H$ components such that $\varphi(a)=-id$). We further split $S$ into subsets $S_{e}$ and $S_{o}$ of almost equal size, where the $\mathbb{Z}_{p}$ component of each element in $S_{e}$ is larger than the $\mathbb{Z}_{p}$ components of all of the elements in $S_{o}$. We then prove the existence of an ordering $\mathbf{x}$ of the elements of $P,N,Z,S_{e},$ and $S_{o}$ such that $\delta,\mathbf{x}$ is a sequencing, where $\delta$ is the prespecified product of the elements of the $D_{j}$s. At the same time, we derive an important bound on the size of overlap between subsets of $\mathbb{Z}_{p}\rtimes_{\varphi}H$ and the set $\delta$ times the partial products of $\mathbf{x}$}.
@@
Finally, we describe a two-step method of ordering the $D_{j}$s. First, we randomly split each $D^{o}_{j}$, each $D^{e}_{j}$ (when $j\in L_{1}$), and each remaining $D_{j}$ (when $j\in L_{0}$) into a four-set partition. Here, we use a probabilistic argument to prove that, given a sequencing $\mathbf{x}$ of $E$, we can obtain a sequencing $\tau_{1},\dots\tau_{u},\mathbf{x}$ of $E$ and of the products $\tau_{i}$ of the partitions of the $D_{j}$s. To conclude, we use again a probabilistic argument to prove that if $\mathbf{x}$ satisfies a technical condition on the intersection between suitable subsets and the set $\delta$ times the partial products of $\mathbf{x}$}, then there exists a sequencing of $E$ and each partition $T_{i}$ of the $D_{j}$s. This gives us Theorem 1.3.
3. Semidirect products
@@
In this section, we consider groups that are semidirect products of the form $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$ for some group homomorphism $\varphi:H\rightarrow Aut(\mathbb{Z}_{p})$. We recall that the group operation is defined as
@@
is an automorphism of $(\mathbb{Z}_{p},+)$. It is well known that the automorphisms of $(\mathbb{Z}_{p},+)$ are the multiplications and hence that there exists $\varphi_{a_{1}}\in\mathbb{Z}_{p}$ such that
@@
Here we let $Dih_{p}$ denote the dihedral group $\mathbb{Z}_{p}\rtimes_{\varphi}\mathbb{Z}_{2}$, where $\varphi(1)(x)=-x$ and $\varphi(0)(x)=x$. In this case, the group operation is defined as
@@
where $\varphi_{0}=1$ and $\varphi_{1}=-1$. If we set $id:\mathbb{Z}_{p}\rightarrow\mathbb{Z}_{p}$ to be the map id(x)=x-id(x)=-x, then dihedral groups are an example of a semidirect product $\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $\varphi:H\rightarrow Aut(\mathbb{Z}_{p})$ satisfies the following property:
$(\star)\qquad\qquad\varphi(h)\in\{id,-id\}.$
In the following, we consider only semidirect products that satisfy property $(\star)$.
Given a not necessarily abelian group $G$, a subset $D=\{d_{1},\ldots,d_{r}\}\subseteq G$ is dissociated if
@@
Let $B\subseteq G$ be a finite subset of a not necessarily abelian group. If $D=\{d_{1},\ldots,d_{r}\}$ is a maximal dissociated subset of $B$, then
@@
$d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots d_{\sigma(i-1)}^{\epsilon_{\sigma(i-1)}}\cdot b\cdot d_{\sigma(i)}^{\epsilon_{\sigma(i)}}\cdots\cdot d_{\sigma(r)}^{\epsilon_{\sigma(r)}}=id.$
@@
$b=\left(d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots d_{\sigma(i-1)}^{\epsilon{\sigma(i-1)}}\right)^{-1}\left(d_{\sigma(i)}^{\epsilon(i)}\cdots\cdot d_{\sigma(r)}^{\epsilon_{\sigma(r)}}\right)^{-1}$
@@
Next, the following lemma relates the size of a set and its dimension.
@@
Let $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $H$ is abelian and $\varphi$ satisfies property $(\star)$. Let $B$ be a finite subset of $G$. Then $|B|\leq 5^{\dim(B)}$.
@@
Let $D=\{(x_{1},a_{1}),(x_{2},a_{2}),\dots,(x_{r},a_{r})\}$ be a maximal dissociated subset of $B$ of size $dim(B)$. Due to the previous lemma, this means that any element $(x,a)$ of $B$ belongs to $\operatorname{span}(D)$ and hence
@@
$(x,a)$ $=\left(\pm x_{\sigma(1)}\pm\varphi_{a_{\sigma(1)}}x_{\sigma(2)}+\dots\pm\varphi_{a_{\sigma(1)}+\cdots+a_{\sigma(t-1)}}x_{\sigma(t)},\sum_{i=1}^{t}\pm a_{\sigma(i)}\right)$
@@
where $(\gamma_{i},\lambda_{i})\in\{(0,0),(1,1),(-1,1),(1,-1),(-1,-1)\}$. Since we have five choices for each pair $(\gamma_{i},\lambda_{i})$, it follows that $|B|\leq 5^{\dim(B)}$. ∎
Following [4], we state a lemma that allows us to “rectify” small sets. We define the parameter
(1) $R=R(A):=c_{1}\max\left((\log p)^{1/2},\frac{\log p}{\log|A|}\right),$
@@
Let $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $H$ is abelian and $\varphi$ satisfies property $(\star)$. If $B\subseteq G$ is a nonempty subset of dimension $\dim(B)<R$, then there is some $\phi\in Aut(G)$ such that the image $\pi_{1}(\phi(B))$ (where $\pi_{1}$ is the projection over the $\mathbb{Z}_{p}$ component) is contained in the interval $(-\frac{p}{100|B|},\frac{p}{100|B|})$.
@@
Given $\lambda\in\mathbb{Z}_{p}$, we consider here the family of maps $\phi_{\lambda}:G\rightarrow G$ defined as
@@
Note that, if $\lambda\not=0$, then $\phi_{\lambda}\in Aut(G)$.
@@
$\{\lambda(x_{1},\dots,x_{r}):\lambda\in\mathbb{Z}_{p}\}\subseteq(\mathbb{Z}_{p})^{r}.$
Due to the pigeonhole principle, there exists distinct $\lambda_{1},\lambda_{2}\in\mathbb{Z}_{p}$ such that $\|\lambda_{1}x_{i}-\lambda_{2}x_{i}\|\leq p^{1-1/r}$ for all $i\in[1,r]$. Set $\lambda\vcentcolon=\lambda_{1}-\lambda_{2}$, we have that $\phi_{\lambda}\in Aut(G)$ and $\pi_{1}(\phi_{\lambda}(D))\subseteq[-p^{1-1/\dim(B)},p^{1-1/\dim(B)}]$. Since $B\subseteq\operatorname{span}(D)$, we have
@@
The thesis follows if we prove that
(2) $\dim(B)p^{1-1/\dim(B)}<p/(100|B|)$
as long as $c_{1}$ is chosen to be sufficiently small.
Set $h=dim(B)$, Equation (2) is equivalent to
$\log({hp^{1-1/h}})=\log(h)+\log{p}-1/h\log{p}<\log{p}-\log(100)-\log{|B|}$
and so to
(3) $h\log(h)+h\log(100)+h\log|B|<\log p.$
We first consider the case where $R=c_{1}(\log p)^{1/2}$. We note that due to Lemma 3.2 this relation is implied by
$h\log{h}+2h\log{10}+h^{2}\log{5}<\log{p}$
which holds since we have assumed that $h<R=c_{1}(\log p)^{1/2}$.
Let know consider the case where $h<R=c_{1}\log p/\log|B|$. Here Equation (3) holds since for $c_{1}$ small enough we have
$1/3\log p\geq\max\left\{h\log h,h\log(100),h\log|B|\right\}.$
Before stating the structure theorem for semidirect products, we prove the following lemma, which allows us to modify a dissociated set by absorbing a given element.
@@
Let $G$ be a not necessarily abelian group and let $D_{1}\cup D_{2}$ be a partition of a dissociated subset $D$ of $G$. For every element $x\in G\setminus\{id\}$, either $D_{1}\cup\{x\}$ or $D_{2}\cup\{x\}$ is dissociated.
@@
Now we show that any subset $A$ of $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$, where $H$ is abelian and $\varphi$ satisfies property $(\star)$, can be split into a family of dissociated sets and a remainder set $E$ of small dimension which can be embedded, as a consequence of the rectification lemma, into $\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)$. We assume that $A$ is not contained in $\{(0,h):h\in H\}$ since $H$ is strongly sequenceable and therefore there exists a sequencing of $A$.
Using standard asymptotic notation, we say that $f\asymp g$ if there exists a constant $C>0$ such that $f=Cg$. We also define $f(p)=o(g(p))$ if $\lim_{p\to\infty}f(p)/g(p)=0$.
@@
Let $p$ be large enough and let $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $H$ is abelian, is strongly sequenceable, and $\varphi$ satisfies property $(\star)$. For every nonempty subset $A\subseteq G\setminus\{id\}$, there is some $\phi\in Aut(G)$ such that $\phi(A)$ can be partitioned as
@@
    Each $D_{j}$ is a dissociated set of size $|D_{j}|\asymp R$ and $|D_{j}|$ is a multiple of eight.
@@
    We can split the interval $1,s]=L_{0}\cup L_{1}$ so that $1$ and $s$ belong to the same set $L_{i}$, and for any $j\in L_{1}$ we have that $D_{j}\subseteq\{(x,a):x\in\mathbb{Z}_{p}\}$ with $\varphi(a)=-id$, and for any $j\in L_{0}$ we have that $D_{j}\subseteq\{(x,b):x\in\mathbb{Z}_{p}\}$ with $\varphi(b)=id$.
@@
    Given $j\in L_{1}$, we split $D_{j}$ into $D_{j}^{o}=\{d_{j,1}^{o},\ldots,d_{j,|D_{j}^{o}|}^{o}\}$ and $D_{j}^{e}=\{d_{j,1}^{e},\ldots,d_{j,|D_{j}^{e}|}^{e}\}$ where $|D_{j}^{o}|=|D_{j}^{e}|=|D_{j}|/2$. Then we set
    $\delta\vcentcolon=\left(\prod_{j\in L_{1}}d_{j,1}^{o}\cdot d_{j,1}^{e}\cdot d_{j,2}^{o}\cdot d_{j,2}^{e}\cdots d_{j,|D_{j}^{o}|}^{o}\cdot d_{j,|D_{j}^{e}|}^{e}\right)\left(\prod_{j\in L_{0}}\prod_{d\in D_{j}}d\right)$
    which does not depend [^2] We will use the notation $\prod_{d^{o}\in D^{o},d^{e}\in D^{e}}d^{o}\cdot d^{e}$ without specifying the orderings if the result does not depend on them. on the orderings on $D_{j}^{o}$, $D_{j}^{e}$, $D_{j}$, and on $L_{0}$ and $L_{1}$. Then $\delta=(z_{0},a_{0})$ where $\varphi(a_{0})=id$, $z_{0}\not=0$, and